I turned on the HEPA at the south end during the LSC. Sorry I ment to turn it off.
Overall a "meh" night for locking I think. The script to all-RF worked several times earlier in the evening, although it was delicate and failed at least 50% of the time. Later in the evening, we couldn't get even ~10% of the lock attempts all the way to RF-only.
Den looked into angular things tonight. With the HEPA bench at the Xend on (which it was found to be), the ETMX oplevs were injecting almost a factor of 10 noise (around 10ish Hz?) into the cavity axis motion (as seen by the trans QPD) as compared to oplevs off. Turning off the HEPA removed this noise injection.
Den retuned the QPD trans loops so that they only push on the ETMs, so that we can turn off the ETM oplevs, and leave the ITMs and their oplevs alone.
We are worried again about REFL55. There is much more light on REFL55 than there is on REFL11 (a 90/10 beam splitter divides the light between them), and we see this in the DC output of the PDs, but there seems to be very little actual signal in REFL55. Den drove a line (in PRCL?) while we had the PRMI locked with the arms held off resonance, and REFL55 saw the line a factor of 1,000 less than REFL 11 or REFL165. The analog whitening gain for REFL11 is +18dB, and for REFL55 is +21dB, so it's not that we have significantly less analog gain (that we think). We need to look into this tomorrow. As of now, we don't think there's much hope for transitioning PRMI to REFL55 without a health checkup.